Circular No. 7995

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SMALL MAGELLANIC CLOUD W. Liller, Viña del Mar, Chile, reports his discovery, on Technical Pan

exposures taken on Oct. 16.20 UT, of a possible nova (mag 12.3) near the center of the SMC in an extremely crowded star field. A CCD exposure on Oct. 17.204 yields broadband V mag 12.66 and position α = 0^{h}56^{m}30^{s}.4 ± 0^{s}.2, δ = −72^{o}36′29′′.4 ± 1′′.1 (equinox 2000.0). No star brighter than mag 14 was seen on a pair of exposures taken on Oct. 9.29.

SUPERNOVA 2002gw IN NGC 922 Independent reports of the discoveries of an apparent supernova in

NGC 922 have been received from Takao Doi (Weimar, TX) and L. A. G. Monard (Pretoria, South Africa) — both with 0.30-m reflectors and unfiltered CCD cameras. Doi provides the following position for SN 2002gw: α = 2^{h}25^{m}02^{s}.97, δ = −24^{o}47′50′′.6 (equinox 2000.0), which is 21′′ west and 33′′ south of the center of NGC 922. Monard gives position end figures 02^{s}.97, 51′′.9. They note that it is not visible on Digitized Sky Survey images — Monard citing limiting red mag 19.5 and Doi noting a faint, elongated (north-south) nebulosity almost at the same position. Magnitude estimates for the new object: Sept. 12.052 UT, [18.5 (Monard); Oct. 13.27, 17.3 (Doi); 15, 17.3 (Doi); 16.027, 17.6 ± 0.3 (Monard); 16.832, 17.6 (Monard).

SUPERNOVA 2002gx Further to IAUC 7990, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 19.0) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Oct. 13.40, 13.41, and 13.42 UT (confirmed on Oct. 16 at mag 19.2). The new object is located at α = 2^{h}31^{m}44^{s}.69, δ = +40^{o}16′39′′.5 (equinox 2000.0), which is 6′′ east and 7′′ south of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on 2001 Dec. 23 (limiting mag 20.2 at S/N = 3).

COMET C/2002 O4 (HÖNIG) Several observers have noted that this comet has lost its condensation

and is fading rapidly. K. Kadota, Ageo, Saitama, Japan (0.30-m reflector) reports (via S. Nakano) m_{1} = 13.3 on a CCD image taken on Oct. 10.82 UT, but Y. Ezaki (Toyonaka, Osaka, Japan, 0.20-m reflector) could not find the comet to mag ∼ 15 on an image taken two nights later.

2002 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green